The time it takes for Jill to grade a set of tests depends on the amont of coffee that she drinks in the morning. After the customary 12 cups of coffee, a set of tests would only take 40 minutes to grade, on a slow morning though (when she only has 4 cups of cofee, a set of tests would take her one hour and twentytwo minutes. Write an equation expressing the grading time in terms of numbers of cups of coffee.
Answer by stanbon(75887)

The time it takes for Jill to grade a set of tests depends on the amont of coffee that she drinks in the morning. After the customary 12 cups of coffee, a set of tests would only take 40 minutes to grade, on a slow morning though (when she only has 4 cups of cofee, a set of tests would take her one hour and twentytwo minutes. Write an equation expressing the grading time in terms of numbers of cups of coffee.
------
You have 2 points relating time and cups: (12,40) and (4,82)
-----
slope = (82-40)/(4-12) = 42/-8 = -21/4
-----
Form: y = mx + b
Solve for "b":
82 = (-21/4)4 + b
b = 103
---
Equation:
y = (-21/4)x + 103
=====================
Cheers,
Stan H.
=====================